### v1

# Workload Estimation for Unknown Tasks: A Survey of Machine Learning Under Distribution Shift

Abstract
Human-robot teams involve humans and robots collaborating to achieve tasks under various environmental conditions. Successful teaming will require robots to adapt autonomously to a human teammate’s internal state. An important element of such adaptation is the ability to estimate the human teammates’ workload in unknown situations. Existing workload models use machine learning to model the relationships between physiological metrics and workload; however, these methods are susceptible to individual differences and are heavily influenced by other factors. These methods cannot generalize to unknown tasks, as they rely on standard machine learning approaches that assume data consists of independent and identically distributed (IID) samples. This assumption does not necessarily hold for estimating workload for new tasks. A survey of non-IID machine learning techniques is presented, where commonly used techniques are evaluated using three criteria: portability, model complexity, and adaptability. These criteria are used to argue which techniques are most applicable for estimating workload for unknown tasks in dynamic, real-time environments.
Index Terms:
human-robot teams, non-iid machine learning, workload estimation for unknown tasks
publicationid: pubid:
I Introduction
Deploying human-robot teams in uncertain environments requires robots to have a dynamic understanding of humans’ internal state. This dynamic understanding must account for real world complexities in order to enable fluid interactions between a robot and a human teammate. Workload represents how hard a person is working, and can be decomposed into workload components (e.g., cognitive, speech, auditory, visual, gross motor, fine motor, and tactile). Incorporating accurate workload models will enhance the robot’s understanding of its human teammate, allowing the robot to manage a human’s workload level, either through intelligent task redistribution or modulation of how the robot interacts with the human. However, real world human-robot teams will inevitably encounter unseen tasks and environments.
An adaptive teaming system requires machine learning models that can estimate workload accurately in both known and unknown situations. Further, workload estimates are susceptible to individual differences across individuals performing the same task, the same individual performing different tasks, and even for the same individual performing the same task on different days. Workload is also influenced by factors, such as experience, stress, and fatigue [1]. Generalizing to novel situations while also accounting for these factors is difficult for traditional machine learning models, as these standard approaches assume data consists of independent and identically distributed (IID) samples [2]. However, this assumption does not hold for estimating workload for new tasks.
Consider first response robotics. Urban firefighters will work alongside robots to perform a wide range of tasks [3, 4]. Firefighters have historically used robots to gather aerial imagery of buildings, but more advanced control algorithms will enable the use of similar robots to extinguish fires in difficult to access locations [4]. This new tasks require different cognitive and visual demands, as well as different degrees of robot control, supervision, and interactions.
Prior work established that standard machine learning methods are capable of generalizing across individuals for similar tasks (e.g., [1, 5, 6]), but there is no widely accepted approach that can generalize to unknown tasks [7]. Generalizing across tasks is difficult primarily due to a shift in the relative balance of workload components; thus, various physiological signals are of particular importance to a specific task and the magnitude of that significance varies between situations [1].
This dynamic relationship between physiological signals and an individual’s underlying workload is exemplar of the variability of human behavior, which makes the development of machine learning models that generalize in the real world difficult. Gathering the necessary data to develop these models exacerbates this problem, as experimental human-subject evaluations rarely capture the fully breadth of human activity within real world domains. Similar complexity is seen across human-robot interaction (HRI) application domains [8, 9, 10]. The real world variability of human behavior and activity violates the IID assumption, requiring non-standard machine learning techniques (e.g., continual learning, meta-learning) to successfully overcome these challenges [11].
The manuscript surveys applicable non-IID machine learning methods. Specifically, the non-IID methods are evaluated to assess their viability for developing models of the human’s internal state (e.g., workload estimation) that can accommodate real world variability. Workload estimation for unknown tasks is an exemplar of the challenge in constructing machine learning models for HRI application domains that are characterized by non-IID data. Section II discusses the fundamental of objective workload estimation and distribution shifts that characterize non-IID machine learning. Section III lays out the notation and evaluation criteria. Section IV evaluates a wide range of non-IID machine learning methods, and Section V discusses the most applicable methods for workload estimation for unknown tasks. Finally, Section VI concludes and presents potential future work.
II Background
Workload is a complex, dynamic, individual-specific, non-linear construct, and developing models that can account for these complexities is a non-trivial task. Fundamentally, estimating workload for unknown tasks is a problem of generalization and recent work has established that existing workload assessment methods encounter difficulties generalizing across tasks, either known or unknown [7] [12]. Modern techniques (e.g., meta-learning) approach machine learning from a different perspective and possess the potential to build more flexible workload estimation models.
Humans have limited resources for processing information, making decisions, and dealing with physical stress. Formally, workload is the degree of activation of a finite pool of resources while performing a task [7] that is influenced by environmental factors, situational factors, and internal characteristics (e.g., fatigue, experience). Workload has been shown to vary across individuals, across tasks, and over time [13] [14]. Workload can be divided into different components based upon the types of resources being utilized: cognitive, speech, auditory, visual, and physical [15], and physical can be further subdivided based upon the nature of physical work: gross motor, fine motor, and tactile [16].
Workload estimation methods either rely on subjective (e.g., NASA Task Load Index [17], In-Situ Surveys [18]) or objective metrics. Subjective metrics fail to provide a continuous measure and are sensitive to the bias inherent to self-reports [19] [20]. Robots operating in dynamic environments need high-frequency workload measures to adequately adjust to their human teammates’ current state. Objective estimation methods use machine learning to learn the relationship between physiological metrics, collected via wearable sensors, and an individual’s underlying workload (e.g., [21, 22, 23, 24, 25]).
Human-robot teams that are deployed in unstructured, dynamic environments cannot rely on environmentally embedded sensors (e.g., built-in cameras), or complex sensor systems (e.g., electroencephalogram (EEG)). Thus, wearable sensors (e.g., a heart-rate monitor) can serve as the primary data source, enabling the human and the robot to act independently. Each workload component corresponds to a different set of physiological metrics, though some metrics may correspond to multiple components. A good example is pupil diameter, which is positively correlated with cognitive and visual workload [16]. Robots need a reliable estimate of the distribution of workload across the components, as tasks will change over time, and differing tasks impact different components.
Prior work established four metrics for evaluating workload estimation algorithms: Sensitivity, Diagnosticity, Suitability, and Generalizability [16]. Sensitivity represents an algorithm’s ability to reliably detect workload levels (i.e., $\geq$ 80% accuracy or $\leq$ 5 root mean squared error (RMSE)). Diagnosticity refers to the capability of detecting different types of workload. Suitability refers to an algorithm’s ability to assess the complete overall workload state. The last metric, generalizability, refers to the algorithm’s ability to generalize across individuals and tasks. This metric defines “generalizing across tasks” as an algorithms capacity to asses the complete overall workload state (i.e., each workload component) and defines “generalizing across individuals” as an algorithms capacity to maintain an accuracy of $\geq$ 80% or an RMSE $\leq$5 for all individuals.
Most existing methods fail to meet these criteria, as they primarily perform binary classification of cognitive workload (e.g., [26, 27, 28, 29]). Many methods are either suitable (e.g., [30, 31, 32, 33]) or sensitive (e.g., [34, 35, 5, 36]), but very few algorithms are diagnostic. Further, little work has been done to evaluate the generalizability of existing workload estimation methods.
Prior work tends to focus on either generalizing across individuals or generalizing across tasks. Many algorithms employed a leave-one-subject-out cross-validation strategy to validate generalization across individuals (e.g., [37, 24, 38, 39]); however, generalizing across tasks, either known or unknown, is rarely evaluated. Algorithms specifically developed to generalize across tasks have received increased attention, and are typically referred to as cross-task workload estimation algorithms [40]. Much of this work evaluated an existing algorithm’s ability to generalize across tasks, but those algorithms did not achieve either a $\geq 80\%$ accuracy, or a $\leq$ 5 RMSE on arbitrary tasks (e.g., [40, 41, 42, 43, 44]). Thus, these algorithms are sensitive, but not generalizable. Other algorithms achieved $\geq 80\%$ accuracy across tasks, but only perform binary classification (e.g., [45, 46, 47]); thus, failing to meet the generalizability criterion.
A support vector machine-based workload estimation algorithm that uses Recursive Feature Elimination to optimize the input features was shown to estimate cognitive workload on arbitrary tasks [48]. This method achieved $<$ 1.0 mean squared error on cross-task workload estimation. Similar methods were developed by Guan et al. [23] and Zhou et al. [49]. Additionally, deep learning techniques (e.g., attention mechanisms) that estimate workload have shown promise [50]. However, the majority of existing methods focus solely on cognitive workload and rely on high data volumes from a static sensor (i.e., EEG [51, 52, 53]). Thus, none of the methods are viable for use in dynamic, uncertain environments.
Real world human-robot teams will inevitably encounter unique scenarios when deployed. An adaptive teaming system requires models that can accurately estimate workload in both known and unknown scenarios, as well as accommodate individual differences [14]. Generalizing to unknown tasks while also accounting for individual differences is difficult for traditional machine learning, as the standard approaches assume data consists of independent and identically distributed (IID) samples [2]. Generalizing across tasks is difficult primarily due to a shift in the relative balance of the workload components associated with a task; thus, various physiological signals are of particular importance to a specific task and the magnitude of that significance varies between situations [1]. This shift in the balance of workload components and physiological signals violates the IID assumption, which requires methods that use non-standard machine learning techniques [2].
III Notation and Evaluation Criteria
Dynamic human-robot teams will rely on machine learning methods to enable a robot to quantify different aspects of their human teammate (e.g., workload). A major challenge is that these teams will continuously encounter new scenarios that the robot’s underlying machine learning methods cannot account for. These new scenarios constitute a meaningful distribution shift from conventional machine learning method’s training dataset (i.e., violate the IID assumption). Specifically, a distribution shift indicates scenarios where training and testing data come from different distributions. Understanding the difference between non-IID machine learning methods is key for building human-robot teams that can handle the variability of real world problem domains.
The workload estimation and the machine learning literature both cover a large spectrum of work. Naturally, there are terms that frequently occur in both fields which require disambiguation. The term task is used in the workload literature to refer to a job an individual may perform (e.g., lifting boxes, mental calculations), but the machine learning literature uses the term to refer to a learning task for a machine learning model. The term domain is often used to characterize the input space for a machine learning problem. For example, performing object detection under various weather conditions (e.g., sunny, cloudy) can be viewed as two domains characterized by the weather [54]. Domain may also refer to a set of activities a human performs for a particular job (e.g., wildland firefighters).
The term learning task will refer to instances of a machine learning problem. The term task, without any adjectives, will refer to a task conducted by a human. The term domain will refer to different input spaces that a machine learning algorithm may operate on. The term application domain will refer to applications where human-robot teams may be deployed.
Definitions of domain and learning task, in the context of distribution shift, have long been established [55]:
Definition III.1
(Domain) A domain $D$ is a two-tuple $(\mathcal{X},P(X))$. $\mathcal{X}$ is the feature space of $D$ and $P(X)$ is the marginal distribution, where $X=\{x_{1},…,x_{n}\}\in\mathcal{X}$.
Definition III.2
(Learning Task) A learning task $L$ is a two-tuple $(\mathcal{Y},f(x))$ associated with a specific domain $D$. $\mathcal{Y}$ is the label space of $D$ and $f(x)$ is an objective predictive function for $D$, which is learned from the training data. $f(x)$ can be written as a conditional probability distribution $P(y|x)$.
Grounding this notation in workload estimation helps to provide an intuitive explanation of the different variables. A particular workload estimation instance represents the domain, $D$, which is characterized by the set of physiological metrics, $\mathcal{X}$. The marginal distribution, $P(X)$, represents the probability distribution over the physiological metrics’ possible values.
A learning task, $L$, is a machine learning problem that is characterized by a label space $\mathcal{Y}$ and a predictive function $f(x)$, corresponding to a single domain $D$. The label space represents a set of possible workload values, and the predictive function produces a workload value, $y_{i}\in\mathcal{Y}$ given a set of physiological metrics, $x_{i}\in\mathcal{X}$. A predictive function can take the form of a conditional probability distribution $P(\mathcal{Y}|\mathcal{X})$, which represents the probability of assigning a workload value to a given set of physiological metrics.
It is important to note that changes to either $D$, $\mathcal{Y}$, or $f(x)$ constitute a new learning tasks. Changes to a domain (i.e., domain shift) are intuitive, as they represent a distribution shift of physiological metrics brought on by performing a task under different working conditions (e.g., weather). Changes to a label space (i.e., label shift) are characterized by a different range of possible workload values or distribution across workload components, typically brought on by performing a new task. Changes to the predictive function $f(x)$ are less intuitive and more easily explained via task recognition.
A typical task recognition problem is defined as training a predictive function to learn the relationship between a set of physiological metrics and a label space (e.g., (1, 2, and 3)), where each integer represents a unique task. A similar task recognition problem can be defined using the same metrics and the same label space, but where each integer represents the task difficulty (i.e., easy, medium, or hard). These two task recognition problems are semantically different, but the underlying machine learning method does not inherently recognizes these differences; therefore, uniqueness of these two learning tasks cannot be fully specified by the domain and label space. The changes in how the resulting predictive functions, $f(x)$, map the physiological metrics to the output values is a key characteristic in differentiating these two learning tasks.
Numerous methods have been applied to the non-IID machine learning problems, each characterized by the number of distributions and distribution shift type (see Fig. 1). Three evaluation criteria are applied to the presented methods: (1) Portability (2) Model Complexity (3) Adaptability.
Portability refers to the number of distributions used to train and test a model [56]. Learning tasks may consist of multiple source domains (e.g., rural hospital data vs. urban hospital data). A straightforward means of combining information from multiple domains is to construct a single joint distribution; however, data in some domains may be of higher quality than others. Weighting high-quality data over low-quality data is appropriate to maximize performance [56]. Additionally, the separation of source domains may be artificial. Prior work established that modeling domains separately can be beneficial for learning inductive biases that improve performance on unknown learning tasks [57]. Portability can have four values: one-to-one, one-to-many, many-to-one, and many-to-many. One-to-one refers to an algorithm that transfers knowledge between one source and one target domain. One-to-many transfers knowledge from one source domain to many target domains, and many-to-one transfers knowledge from multiple source domains to a single target domain. Many-to-many knowledge transfer occurs between multiple sources to multiple target domains simultaneously.
Figure 1: A high level overview of non-IID machine learning methods. Adapted from Figure 19.11 of [2].
Model Complexity refers to the number of parameters an underlying machine learning model needs for the algorithm to be successful [58]. Machine learning models with more parameters require higher volumes of training data. Early statisticians built simple linear models using hundreds of examples [58], whereas modern deep learning architectures may require tens of billions of examples [59]. Model complexity shares a relationship with both the underlying machine learning model type and problem complexity, as more complex models are often required to solve more complex problems. Workload estimation and many HRI problems are typically conducted in low-data regimes, as they require human subject evaluations for machine learning model development; thus, smaller models are preferred. An example of low model complexity is a random forest model that contains a few hundred decision trees, where the depth of each tree is dataset-dependent; thus, these models typically contain on the order of thousands of parameters [60]. An example of medium model complexity is shallow neural networks (i.e., fully connected neural networks) that typically contain hundreds of thousands of parameters [22], but may contain on the order of millions of parameters. An example of high model complexity is deep neural networks (e.g., convolutional networks, long short-term memory networks (LSTM)), also referred to as deep learning architectures. Deep learning architectures can contain anywhere from tens of millions [61] to over 100 billion parameters [59].
The last metric, Adaptability, refers to the amount of data required to update a machine learning model to solve new problems [62] [63]. Many non-IID machine learning methods seek to account for distribution shifts by updating existing machine learning models’ parameters so that the model can solve new learning tasks. Adaptability is the number of test data points necessary for the updated model to make accurate predictions for new learning tasks, where the goal is to minimize the required data. A machine learning method has no adaptability if it requires the same amount of training data in the source and target domains. Prior work suggested that 5000 examples per class is a sufficient rule of thumb for training deep learning architectures from scratch to achieve a minimum acceptable performance [58]. Therefore, an algorithm is considered to have low adaptability if it requires $>$ 2500 examples per class (i.e., many) to optimize a function in a target domain [63]. A 50% reduction in data required for updating a model is non-trivial, and may still constitute a large dataset. An algorithm has high adaptability if it requires $<$ 10 (i.e., a few) examples to update a model, a small dataset in all contexts [62]. It is important to note adaptability is application domain-specific, as the learning tasks are fundamentally different.
IV Non-IID Machine Learning Methods
There are several non-IID machine learning methods that can account for different distribution shift types, including: transfer learning, domain adaptation, test-time training, continual learning, domain generalization, and few-shot learning. Each method corresponds to a specific distribution shift type and the number of distributions considered (see Figure 1). There are two primary distribution shifts: (1) Domain shift, and (2) Label shift. A domain shift indicates a distribution shift of the input features, and label shift indicates a shift of the output values. Label shifts can be characterized by either a new distribution of output values, or a new set of values.
Standard IID methods only account for data from a single distribution, which can lead to performance degradation in the presence of distribution shifts [2]. Non-IID methods seek to account for these shifts by intelligently incorporating information from multiple distributions (i.e., domains). Further, these methods either account for exactly two distributions or from more than two distributions, where the boundaries between distributions may be real or artificial. The variability of human behavior is most aptly described by learning tasks with multiple distribution shifts across domains, as humans perform diverse tasks in different ways in a wide range of environments. However, the data can be decomposed or re-organized such that any of the presented techniques can be applied. Data from a single, noisy distribution may be decomposed into multiple distributions to improve generalization to new tasks [64], or data from multiple distributions may be aggregated into a joint distribution to increase data volume and use more complex models [63].
Many of these methods have been successfully applied to task recognition using wearable sensors (e.g., [65, 66, 67, 68]), which shares core commonalities with workload estimation. Thus, these methods provide insight into the capabilities of the underlying machine learning models to account for distribution shifts in relevant application domains. Prior work has primarily focused on evaluating standard IID methods’ ability to generalize to new situations, where many methods fail to generalize to new tasks (e.g., [40, 43, 48, 47]). Relying on standard IID methods ignores the distribution shifts of the physiological signals and workload components as tasks change. No prior work modeled workload estimation for unknown tasks as a non-IID problem. An appropriate non-IID machine learning method for the workload estimation for unknown tasks must be identified, so that the resulting workload estimation algorithm can account for distribution shifts across tasks.
IV-A Domain Adaptation
Domain adaptation considers the setting where training data and test data are characterized by the same features, but are sampled from different distributions.
Definition IV.1
(Domain Adaptation [69]) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, domain adaptation aims to improve the target predictive function $f_{T}(x)$, where $D_{S}\neq D_{T}$.
This distribution shift type is called a domain shift, or a covariate shift. For example, consider multiple datasets with images of the same objects [70]. Changes in the images’ lighting, background, or orientation represent meaningful difference in the input features and are prototypical examples of a domain shift. Domain adaptation is a non-IID machine learning method that seeks to minimize performance degradation in the presence of domain shift. It is important to note that domain adaptation techniques focus on distribution shifts between one source and one target domain (i.e., two distributions), as shown in Figure 1.
Domain adaptation has also been called transductive transfer learning, as many techniques require the presence of source and target domain data during training in order to perform well [55]. Further, there are two forms of domain adaptation: supervised [70] and unsupervised [61]. Supervised domain requires that labels be available for $L_{T}$, and there are no labels available for $L_{T}$ for unsupervised domain adaptation. There tends to be significantly more data available in source domains than the target domains, and using labeled data from both domains typically results in overfitting to the source distribution [70]. Thus, the majority of prior work has focused on unsupervised domain adaptation.
Domain adaptation has been applied in broad range of HRI application domains, including task recognition (e.g., [71, 72, 65, 73, 74]), workload estimation (e.g., [49, 75, 76]), and emotion recognition (e.g., [77, 78, 79, 80]). The vast majority of these applications use adversarial learning techniques [66, 75, 80], though self-supervised and feature representation learning strategies have also been moderately successful [77].
Adversarial domain adaptation consists of three model types: (1) feature extractor, (2) domain discriminator, and (3) label predictor. Seminal work integrated adversarial learning and domain adaptation as a mini-max game [81]. This work demonstrated that domain adaptation can be achieved by jointly optimizing a feature extraction model to maximize the difference of latent features from source and target domain data, via a domain classifier, and to minimize the error in predicted values for learning tasks, via a label predictor. This joint optimization enables the effective learning of a useful latent feature representation. Prior work demonstrated that replacing the discriminative feature extraction models with generative feature extraction models is also an effective means of domain adaptation (e.g., [82, 83, 84]).
Most domain adaptation techniques exhibit one-to-one portability [84], but multi-target domain adaptation (i.e., one-to-many portability) is an active research area [85]. Domain adaptation techniques primarily rely on deep neural networks [49, 75, 77], consisting of millions of parameters; thus, exhibiting a high model complexity. The adaptability of domain adaptation varies. Some approaches require one example in the target domain [86], while others require thousands [66]. Further, access to source and target data during training is a hard requirement for many domain adaptation techniques, which makes them impractical for real world HRI. One-to-one portability, high model complexity, and low adaptability are undesirable properties; thus, domain adaptation techniques are not viable for workload estimation of unknown tasks.
IV-B Transfer Learning
Transfer learning aims to improve the performance of a machine learning model for a target domain by transferring knowledge from a different, but related source domain.
Definition IV.2
(Transfer Learning [63]) Given a source domain $D_{S}$, and a learning task $L_{S}$, a target domain $D_{T}$, and learning task $L_{T}$, transfer learning aims to improve the learning of a target predictive function $f_{T}(x)$ in $D_{T}$, using the knowledge in $D_{S}$ and $L_{S}$, where $D_{S}\neq D_{T}$ and $L_{S}\neq L_{T}$.
Transfer learning is characterized a label shift [87], which can refer to a shift in the relative distribution of output values (e.g., disease prevalence in a rural hospital vs. urban hospital), or a complete change in the labels available. Transfer learning techniques are capable of transferring knowledge from a model trained to identify everyday objects (e.g., cats, planes) in images to a model performing medical image analysis (e.g., ultrasounds, x-rays) [88]. Similar to domain adaptation, transfer learning has been widely applied to HRI (e.g., human task recognition [89, 90], emotion recognition [91, 92, 93, 94]). There are many transfer learning techniques, which can be categorized based on how knowledge is transferred: instance-based (e.g., [95, 96, 97]), feature-based (e.g., [98, 99, 100]), and parameter-based (e.g., [101, 87, 68, 102]). Instanced and feature-based techniques have had moderate levels of success, but are seldom used in modern applications when compared to parameter-based techniques [103].
Parameter-based methods have two phases: pre-training and fine-tuning. Pre-training trains a machine learning model on data from the source domain, $D_{S}$, to solve the source learning task $L_{S}$. Fine-tuning trains the same model on data from the target domain, $D_{T}$, but only updates a subset of the parameters. This two-step training procedure enables the network to learn a useful latent representation during pre-training, and a task-specific transformation of that representation during fine-tuning. This approach has been successful in computer vision [87], robotics [104], and task recognition (e.g., [105, 106, 107, 108]).
Many of these methods use deep learning architectures to transfer knowledge across source/target domain pairs, requiring thousands of examples to be successful. Further, there is evidence to suggest that the model complexity is inversely related to both portability and adaptability [63] [59]. Zhuang et al. evaluated a range of deep learning architectures on several tasks, all of which exhibited one-to-one portability, high model complexity, and low adaptability [63]. Similar findings exist in transfer learning for human activity recognition [105]. Similar techniques were also developed to solve a wide range of natural language problems based on only a few examples [59]. This high adaptability is in direct contrast to the previously discussed transfer learning techniques, where the primary difference is the model complexity. A transformer deep learning architecture was trained to predict the next word in a sentence [59]. This network was the foundation of several downstream networks that performed tasks, including machine translation, question answering, and arithmetic word problem-solving. The architecture consisted of 175 billion parameters, requiring significantly more data to train parameter-based techniques that rely on smaller networks (i.e., tens of billions of data points). These results suggest that a higher model complexity can directly improve portability and adaptability.
A core limitation of workload estimation is data volume, as gathering large volumes of wearable sensor data is logistically impractical; thus, model complexity is constrained. It is also desirable that workload estimation algorithms exhibit many-to-many portability, as models may need information across multiple individuals and tasks. Thus, transfer learning is not a viable option for workload estimation of unknown tasks.
IV-C Test-time Adaptation
Domain adaptation techniques learn to anticipate domain shifts by simultaneously training on $D_{S}$ and $D_{T}$; however, target domain data may not be available during training. Test-time adaptation techniques seek to overcome this limitation.
Definition IV.3
(Test-time Adaptation [109]) Given a source domain $D_{S}$ and a learning task $L_{S}$, as well as a target domain $D_{T}$ and learning task $L_{T}$, test-time adaptation aims to improve the learning of a target predictive function $f_{T}(x)$ based on an existing predictive function $f_{S}(x)$, where $D_{S}\neq D_{T}$.
Test-time adaptation techniques seek to overcome domain shifts by separating the training process into two distinct phases: (1) training on $D_{S}$ data, and (2) updating that model with $D_{T}$ data. This separation enables the resulting model to adapt at test-time by allowing for the continuous updating of model parameters, as shown in Figure 1 [109].
Prior work applied test-time adaptation techniques to egocentric task recognition [110] and adaptive policy optimization for assistive robots [111]. However, test-time adaptation techniques have not been widely applied to HRI application domains. Nevertheless, these techniques hold promise because they overcome the limitation of training on source and target domain data simultaneously. These techniques can be categorized into as either self-supervised, or entropy-based.
Self-supervised techniques use a dual-output neural network and a two-phase training procedure [109]. Labeled data from the $D_{S}$ undergoes some deterministic transformation (e.g., image rotation). The neural network is trained on this transformed data to predict both the original label (e.g., image class) and the corresponding transformation (e.g., rotation angle). The unlabelled data from $D_{T}$ undergoes the same deterministic transformation. The shared feature extraction model’s parameters are updated based on the output of the self-supervised proxy learning tasks. This two-phase procedure decouples training from adaptation and enables the model to adapt at test-time. Many techniques take inspiration from this work (e.g., [112, 113, 114]); however, the development of an application domain relevant proxy task is non-trivial, limiting the generalizability of self-supervised techniques.
Entropy-based techniques eliminate the need for a self-supervised proxy task by using a batch processing adaptation phase (e.g., [115, 116, 117]). Batches of data from $D_{T}$ are passed through the shared feature extractor model, and the Shannon Entropy [118] is calculated for each batch. The adaptation phase updates the shared feature extractor to minimize the average entropy over all batches [119].
Fundamentally, test-time adaptation is domain adaptation without access to target domain data during training. Therefore, test-time adaptation exhibits one-to-one portability, high model complexity, and low adaptability. It is difficult to definitively state whether these techniques are applicable to workload estimation for unknown tasks, as the relationship between model complexity and adaptability for these techniques is unknown. Prior work has primarily focused on deep learning architectures, but these techniques are general enough to apply to shallow neural networks. It is possible that smaller networks may possess sufficient adaptability to account for real world variability, but the efficacy of test-time adaptation on these networks has not been thoroughly evaluated.
IV-D Continual Learning
Continual Learning techniques train a model to perform a sequence of non-IID learning tasks, where each task is characterized by a unique domain shift (see Figure 1).
Definition IV.4
(Continual learning [120]) Given a sequence of domains $D$ = $D_{s_{1}},...,D_{s_{n}}$, where $n>0$, and a sequence of learning tasks $T$ = $L_{s_{1}},…,L_{s_{n}}$, where learning task $L_{s_{i}}$ corresponds to domain $D_{s_{i}}$, continual learning aims to optimize the predictive function $f_{i}(x)$ using knowledge from previously seen domains $D_{j}\in D$, where $i\leq j$.
These techniques gradually acquire knowledge to optimize performance on the latest learning task, and some techniques add the constraint of avoiding performance degradation on all prior tasks [121]. Continual learning is also referred to as lifelong learning, sequential learning, or online learning [122]. There is a wide range of techniques, including knowledge distillation (e.g., [123, 124, 125]), memory-based methods (e.g., [126, 127]), and parameter isolation (e.g., [128, 129]).
Continual learning techniques are one of the more popular techniques within HRI, having been successfully applied to problems of human task recognition using wearable sensors (e.g., [130, 131, 132, 133]), affective computing (e.g., [8, 9, 134]), and adaptive human-robot interactions (e.g., [8, 135, 136]).
Conventional continual learning assumes the new learning tasks arrive one at a time, and the data in the domains is stationary [137]. Thus, continual learning techniques can be trained offline as the number of tasks is known a priori. These techniques are referred to as multi-head techniques, as separate models are trained for each learning task [138]. However, knowing all the tasks ahead of time is not always realistic. Other continual learning techniques make no assumptions about the distribution of learning tasks, and seek to identify the task online [138]. These techniques are referred to as single-head techniques, since a single model is trained for all tasks.
Learning tasks within a sequence can either be independent or dependent. Independent learning tasks are separate and may not share similarities. Dependent learning tasks represent a sub-sequence of learning tasks that must be executed in a particular order, where learning task $L_{i}$ holds key information about learning task $L_{i+1}$. Many of these techniques rely heavily on deep neural networks, similar to both domain adaptation and transfer learning techniques [139, 140].
The model complexity of continual learning techniques varies; however, there is evidence that model complexity is not directly tied to performance, as the machine learning model choice is problem specific (e.g., Gaussian processes [136], shallow neural networks [127], and deep networks [122]). Many techniques exhibit low adaptability, requiring thousands of examples to perform well [122]. Portability is challenging to define for continual learning, as it exhibits many-to-many portability, but does so only in a sequential fashion. Many-to-many portability is desirable, but the sequential nature is an unnecessary constraint. Additionally, some continual learning methods only optimize performance on the latest learning tasks, while other optimize performance on all prior learning tasks. Portability and adaptability are key qualities required to estimating workload successfully for unknown tasks; thus, continual learning is a subpar option.
Figure 2: Comparison of applying the technique Model Agnostic Meta-learning [141] to: (a) Domain Generalization, sub-problems experience a domain shift, and (b) Few-shot Learning, sub-problems experience a domain and label shift.
IV-E Domain Generalization
Domain generalization considers the setting where training data from multiple source domains is characterized by the same features that are sampled from different distributions [142]. Specifically, domain generalization is the process of transferring knowledge acquired from several related source domains and applying it to previously unknown domains [143], as indicated in Figure 1.
Definition IV.5
(Domain Generalization [142]) Given a set of source domains $DS$ = $D_{s_{1}},…,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source tasks $LS$ = $L_{s_{1}},…,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target task $L_{t}$ that corresponds to $D_{t}$, domain generalization aims to optimizes a target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$.
The goal is to leverage unique cross-domain information to improve the machine learning model’s generalization. These methods assume the construction of a joint distribution of source domains aggregates away key information, and that the varying quality and relevance of data across source domains helps inform a model in unknown target domains. A key feature of domain generalization is the lack of labeled target domain samples. Domain generalization techniques can be broadly classified into three categories: domain-invariant representation learning (e.g., [144, 145, 146]), data manipulation (e.g., [147, 148, 149]), and learning strategy (e.g., [150, 151, 152]).
The primary technique for domain-invariant representation learning is domain alignment, which seeks to minimize the difference across source domains, such that a latent representation consisting of only essential features is learned [153]. Domain alignment techniques measure this difference in many ways, including minimizing moments [154], contrastive loss [155], KL divergence [156], or maximum mean discrepancy [157]. Prior work combined domain alignment metrics with domain-adversarial learning [157]. The primary difference between domain-adversarial techniques for domain generalization and domain adaptation is that the source/target discriminator must differentiate between multiple source domains.
Data manipulation techniques seek to artificially enhance the size and diversity of a dataset via data augmentation or generation. Data augmentation performs deterministic transformations over existing data to create new examples [158]. However, data augmentation is not domain-agnostic, as useful transformations are specific to an application domain. Data generation uses generative machine learning models (e.g., Variational Autoencoders [147], Generative Adversarial Networks [148]) to create completely new training examples. These techniques are domain-agnostic, but the underlying models are hard to train and require a high training data volume [159]. These techniques cannot be applied to workload estimation, as generating realistic physiological metrics that correctly correspond to the workload values is non-trivial.
Learning strategy-based domain generalization techniques seek to improve cross domain generalization through novel training methodologies, including: 1) Self-supervsied learning 2) Ensemble learning, and 3) Meta-learning. Self-supervised learning techniques for domain generalization do not differ substantially from self-supervised techniques for test-time adaptation (see Section III.C). Ensemble learning is the process of training multiple machine learning models with varying weight initializations or training data splits, then aggregating the models’ outputs to make predictions. These extensively studied techniques are notoriously computationally expensive [160] and are not viable solutions for real-time HRI domains.
Meta-learning trains a machine learning model on data across source domains, such that that model can be efficiently updated to operate in arbitrary target domains [141]. Broadly, meta-learning is ‘learning to learn’ [161]. Domain generalization typically leverages parameter-based meta-learning techniques (e.g., Model Agnostic Meta-Learning (MAML) [141]) that seeks to construct a neural network that can be adapted based on a few examples (see Figure 2).
A meta-model, parameterized by $\theta$, is trained to predict an initialization for the parameters of another model, $\phi_{i}$, that performs the learning task $L_{i}$. Both $\theta$ and $\phi_{i}$ typically represent a neural network’s weights, but can also represent a subset of the network’s weights, or the another model’s parameters. Each source domain $D_{i}$ and learning task $L_{i}$ is deconstructed into a collection of mini-datasets. These mini-datasets are defined by the support set, $S_{i}$, and the query set, $Q_{i}$, which are small training and test sets, respectively. The support set consists of a few (e.g., 1, 5, 10) examples that the meta-model uses to predict the labels for the query set. The task-specific model’s error is twice differentiated and backpropagated through both models to update $\theta$. This problem formulation constructs the learning process, such that the model learns how to adapt its parameters based on a few examples. This technique is popular in meta-learning [161], and is the foundation for many techniques (e.g, [162, 163, 164, 165, 166]). An alternative approach learns the parameters of an optimizer. An LSTM was trained to learn the exact optimization process of a neural network, such that the new network converged using only a few samples [167]. The technique Meta-Stochastic Gradient Descent combined the LSTM approach and MAML to learn an initialization of the parameters, and an optimal optimizer [168].
The majority of domain generalization techniques exhibit many-to-many portability by construction. Data manipulation is characterized by high model complexity, due to its reliance on generative deep learning architectures [169]. Additionally, these techniques exhibit a range of adaptability, where some instances require several thousands of examples for adaptation [170], and others only require a few examples [171]. MAML and other meta-learning techniques exhibit high adaptability and are general enough to apply to shallow neural networks; thus, exhibiting a low model complexity. Therefore, meta-learning based domain generalization techniques may be applicable to workload estimation for unknown tasks; however, these techniques have been shown to be computationally expensive at both training and inference times [162].
IV-F Few-shot Learning
Few-shot learning aims to train a machine learning model using data across multiple source domains, such that the model can be efficiently adapted to make predictions in some unknown target domain [62].
Definition IV.6
(Few-shot Learning [62]) Given a set of source domains $DS$ = $D_{s_{1}},…,D_{s_{n}}$ where $n>0$, a target domain, $D_{t}$, a set of source learning tasks $LS$ = $L_{s_{1}},…,L_{s_{n}}$ where $L_{s_{i}}\in LS$ corresponds with $D_{s_{i}}\in DS$, and a target learning task $L_{t}$ that corresponds to $D_{t}$, few-shot learning optimizes the target predictive function $f_{t}(x)$ in $D_{t}$, where $D_{t}\notin DS$ and $L_{t}\notin LS$.
Few-shot learning is often characterized by both a shift in the input distribution (i.e., domain shift) and a shift in the target domain labels (i.e., label shift), as indicated in Figure 1. Prior work primarily focused on applying meta-learning techniques to solve few-shot learning problems. These techniques overlap significantly with meta-learning for domain generalization. Meta-learning constitutes the vast majority of popular few-shot learning algorithms (e.g., MAML [141], Prototypical network [172], Conditional Neural Processes [173])). Meta-learning techniques broadly fall into three categories: 1) parameter-based 2) metric-based 3) probabilitstic. Parameter-based meta-learning algorithms were discussed in Section III.G. An illustrative example demonstrating the primary difference between meta-learning for domain generalization and few-shot learning is provided in Figure 2.
Metric-based meta-learning attempts to learn a robust latent representation across source domains, such that data points can be directly compared [64, 174, 175]. Two neural networks are trained to encode data, from the source and target domain, into a latent feature space. One neural network, $g_{\theta}$, encodes the support set (i.e., source domain data), and a second, $f_{\theta}$, encodes the query set (i.e., target domain data). Typically, multiple support set data points are passed through the networks simultaneously to enable the calculation of a range of complex distance metrics.
(a) Conditional Neural Process
(a) Conditional Neural Process
(b) Latent Neural Process
Figure 3: A comparison of two basic neural process types: (a) conditional neural processes and (b) latent neural process. Green boxes represent input and output values, blue boxes represent a neural network, yellow boxes represent aggregated latent features, and the pink circle represent an aggregation function. This figure is adapted from Figure 3 in [176].
Matching networks perform a pairwise comparison between latent features using cosine similarity [64]. Prototypical networks extend this technique by using additional data points to calculate the centroid of each class’ location within the latent space [172]. Each class is represented by $N$ examples, and when $N$ is equal to one, prototypical networks are identical to matching networks [172]. Relation networks decompose the problem into separate modules: an embedding module and a relation module [177]. The embedding module is a neural network designed for feature extraction and the relation module is a non-linear classifier trained on the latent feature representation. Adaptive subspace networks use singular-value decomposition on the latent features to construct the basis vectors of subspaces for each class [178]. This technique maximizes the distance between subspaces by using a Grassmannian geometry-based objective function. Recent work modeled encoded data points probabilistically to improve information contained within latent features (e.g, [179, 180, 181]).
Generally, probabilistic meta-learning typically can be viewed as the combination of a stochastic process (i.e., Gaussian Process [182])) and a neural network. Combining these two different machine learning models takes the best of both worlds by leverage the Gaussian processes’ Bayesian framework and the neural network’s ability to extract latent features from large datasets [176]. Probabilistic meta-learning techniques fall into two broad categories: Neural Processes [183], and Deep Kernel Learning [184].
Neural processes pass the data through a neural network, parameterized by $\theta$, to learn a latent representation. Each support set data point is passed through an aggregation function, $a$, to calculate the latent feature, $z$. This aggregation is performed to maintain indifference under the permutation of latent features, which helps maintain a key theoretical property (i.e., the consistency condition) of stochastic processes [176]). Another neural network, parameterized by $\phi$, uses $z$ and one or more data points (e.g., one-shot vs. few-shot) from the target domain to learn the parameters of a Gaussian Process. There are two main neural process types: latent neural processes, and conditional neural processes, as shown in Figure 3.
Latent neural processes construct a stochastic variable using latent features, which typically takes the form of a Gaussian distribution parameterized by $\mu$ and $\sigma$ [183]. However, by modeling these correlations the underlying predictive model becomes intractable and must be approximated (e.g., Monte Carlo Sampling [185], Variational Inference [183]). Conditional neural processes do not construct a stochastic variable using latent features [173], instead they make the assumption the stochastic process can be fully described by these encoded values (i.e., factorized Gaussian distribution [176]). This assumption makes the underlying predictive model analytically tractable, but ignores any correlations that may exist between latent features. Several neural network architectural components have been incorporated into conditional neural processes with varying degrees of success (e.g., [186, 187, 188, 189, 190, 191, 192, 193]).
Deep kernel learning techniques takes an alternative by constructing an end-to-end model that simultaneously trains a neural network and a Gaussian process, effectively training the network to learn a custom kernel function for that Gaussian process [184]. Deep kernel transfer first applied this technique to few-shot learning by maximizing the marginal likelihood across multiple tasks [194]. This technique learns over a large number of small, but related, examples that train the underlying machine learning model to approximate a prior distribution to transfer knowledge between learning tasks. Deep kernel transfer quantifies uncertainty by maintaining a distribution over parameters of the neural network, which can become computationally expensive for larger networks. Pólya-Gamma augmentation and softmax approximation were explored to overcome this limitation by maintaining a distribution over functions [195]. Deep mean function learning has also been applied as a meta-learning technique, where a neural network is used to learn a custom mean function for a Gaussian process [196]. These and similar techniques have received increased attention, as neural networks have been combined with a range of probabilistic models (e.g., [156, 197, 198, 199, 200]). However, recent results demonstrate that the original deep kernel learning algorithm tends to produce unreliable uncertainty estimates in some practical application domains [201]. Incorporating modern deep kernel learning techniques and practices has the potential to improve the generalization of the presented meta-learning techniques (e.g., [201][202])
Few-shot learning techniques exhibit many-to-many portability and high adaptability. Most few-shot learning techniques were developed using computer vision benchmarks (e.g., [203, 172, 64]), which requires deep learning architectures. However, prior work demonstrated that few-shot learning techniques’ success is directly tied to a good feature extraction model [11]. Therefore, it stands to reason that model complexity is directly related to the learning task’s complexity. Meta-learning techniques have been proven to be successful solutions to both few-shot learning and domain generalization. Further, these techniques exhibit desirable properties for all three presented criteria. Therefore, meta-learning presents a promising option for develop workload estimation for unknown tasks solutions and merit future investigation.
V Discussion
Workload estimation for unknown tasks relies on noisy and varied metrics, as different physiological signals and workload components are of particular importance to each task. Understanding the distribution across workload components is critical to adapting interactions and task allocations to maximize the performance of a human-robot team. Quantifying uncertainty helps enumerate this distribution by contextualizing component-specific workload estimates, as components may rely on overlapping metrics (e.g., fine-motor and tactile).
Domain adaptation and transfer learning are popular methods, but require high model complexity to achieve desired performance levels. These methods are also best suited to one-to-one portability problems with low adaptability. Specifically, these techniques are most applicable to problems where knowledge can be transferred between data-rich source and data-sparse target domains and where all data is gathered prior to training. These techniques do not adapt machine learning models in the moment; thus, are not viable for performing workload estimation for unknown tasks in real world application domains. Test-time adaptation overcomes this particular constraint, but still exhibits one-to-one portability and low adaptability. Existing test-time adaptation techniques applied to smaller neural networks may exhibit higher adaptability, but further investigation is required.
Continual learning has been frequently proposed as a potential solution to non-IID machine learning problems within HRI, but are not applicable to workload estimation for unknown tasks due to their poor portability. Incorporating information for an indeterminant number of tasks, while avoiding performance degradation on all prior tasks are additional constraints that likely make the machine learning problem unnecessarily difficult. These techniques learn how to transform the model between task pairs, whereas other techniques (e.g., domain generalization, few-shot learning) learn to transform from a common prior. Consider an application domain with $N$ tasks. Continual learning must learn $N!$ task-pair independent updates, as the tasks’ order influences the learning process. Domain generalization and few-shot learning techniques only learn $N$ updates, as they can trace each task to a single common prior. These additional portability-related constraints make continual learning a subpar option.
Domain generalization and few-shot learning exhibit desirable properties for all three criteria: high adaptability, one-to-one portability, and low model complexity. These techniques are explicitly designed to update a machine learning model based upon limited and highly variable data, and are flexible enough change the model complexity based upon the learning task’s complexity. Multiple domain shifts from a common prior is the most directly related metaphor for the variability seen in human behavior; thus, domain generalization is the most suited for scenarios where there is no label shift (e.g., workload estimation). Few-shot learning may be more applicable to scenarios where labels change between learning tasks or application domains (i.e., task recognition, affect recognition). The presented meta-learning techniques are general enough to apply to either machine learning problem.
Utilizing individual workload components (i.e., cognitive, speech, auditory, visual, gross motor, fine motor, tactile) to produce measures of overall workload is fundamentally a regression problem [1, 204, 21]. Only a limited number of techniques (e.g., probablistic meta-learning) have been applied to regression problems with those applications being focused on toy problems and contrived benchmarks. Probablistic meta-learning techniques (e.g., deep kernel transfer, conditional neural process) are particularly promising for solving workload estimation for unknown tasks, as producing continuous values for each workload component is a key requirement for informing an adaptive teaming system.
A key research question is if existing workload estimation datasets capture the non-IID nature of human behavior, such that probabilistic meta-learning techniques can accurately estimate workload for unknown tasks. It is likely that new human subjects evaluations are required to generate the necessary data, but designing evaluations that appropriately capture the variability of human behavior is non-trivial. These evaluations must incorporate diverse, but ecologically valid tasks characteristic of the real world, while also considering the degree of difference between IID and non-IID tasks that these techniques can handle. Understanding how probablistic meta-learning techniques can estimate workload accurately for similar unknown tasks, as well as dissimilar unknown tasks is crucial to understanding their real world viability. Additionally, analyzing how the differences between tasks, human-robot interactions, robot capabilities further informs how these techniques can be successfully deployed. Regardless, these techniques meet all three criteria and are the most promising path forward for solving workload estimation for unknown tasks; thus, merit further investigation.
VI Conclusion
Developing models that can estimate the human’s internal state is critical to developing dynamic human-robot teams that can operate in uncertain, dynamic environments. The variability of these environments is characterized by varying human-robot teaming dynamics, diverse tasks executions, and different operating conditions for the team itself. A robot seeking to adapt its behavior to best assist its human teammate must accurate the human’s internal state accurately in any of these conditions; however, real-world variability often violates the IID assumption making standard machine learning methods infeasible. Non-IID machine learning, particular domain generalization and few-shot learning, hold promise in developing models that can account for these varying dynamics.
Acknowledgments
The graduate students have been supported by National Aeronautics and Space Administration through University Space Research Association aware 904186092. The contents are those of the authors and do not represent the official views of, nor an endorsement, by the National Aeronautics and Space Administration.
The presented work was partially supported by ONR grants N00024-20-F-8705 and N00014-21-1-2052. The views, opinions, and findings expressed are those of the authors and are not to be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.